# Pairs of Clark Unitary Operators on the Bidisk and their Taylor Joint Spectra

### v1

@@ 3.3. Taylor Joint Spectrum
While there are several different notions of spectrum for tuples of commuting bounded operators, the Taylor joint spectrum has proven to be a useful tool in the study of shift operators on Hilbert spaces of analytic functions; to mention a few instances of this, see [18, 23, 7] and their lists of references. Comprehensive treatments of the Taylor joint spectrum for tuples of operators can be found in [36, 19] or [30, Chapter IV]. These references also include an overview of other spectral notions (the Harte spectrum, etc.) and conditions for when they coincide. Since we deal exclusively with pairs of operators, we only provide the relevant definitions in that setting.
@@ 4. Clark Unitary Operators $(U_{\alpha}^{1},U_{\alpha}^{2})$
Let $\phi$ be an inner function in $\mathbb{D}^{2}$. We begin by computing the adjoint of the associated Clark embedding $J_{\alpha}$ which was defined in Subsection 3.2. The following result is a direct analog of Clark’s formula for the one-variable adjoint, see e.g., [22, Theorem 11.6].
@@
$\vartheta^{1}_{\alpha}(z):=\bar{\alpha}\frac{(B_{1}\phi)(z)}{1-\bar{\alpha}\phi(0,z_{2})}.$ (4.2)
@@
Now we will use Proposition 4.2 to connect $M_{z_{1}}$ on $L^{2}(\mu_{\alpha})$ with the compressed shift $S_{\phi}^{1}$ on $K_{\phi}$ for certain $\phi$. To do that, recall that $H^{2}_{2}$ is the Hardy space on the disk with independent variable $z_{2}$. If $f\in H^{2}(\mathbb{D}^{2})$, then
@@
$U^{1}_{\alpha}=S_{\phi}^{1}+P_{\phi}P_{H^{2}_{2}}M_{\overline{\vartheta^{1}_{\alpha}}}.$ (4.3)
@@
Let $U_{\alpha}^{1}=(V_{\alpha}^{1})^{*}$. By Proposition 4.2 and the assumption that $J_{\alpha}$ is a surjective isometry (and hence unitary), we have $M_{z_{1}}=J_{\alpha}U^{1}_{\alpha}J^{*}_{\alpha}$ or equivalently, $U_{\alpha}^{1}=J_{\alpha}^{*}M_{z_{1}}J_{\alpha}$. Since $M_{z_{1}}$ is unitary on $L^{2}(\sigma_{\alpha})$, the operator $U_{\alpha}^{1}$ is unitary on $K_{\phi}$. Furthermore, since $\vartheta^{1}_{\alpha}$ is bounded, multiplication by $\vartheta^{1}_{\alpha}$ and by $\overline{\vartheta^{1}_{\alpha}}$ (denoted $M_{\vartheta^{1}_{\alpha}}$ and $M_{\overline{\vartheta^{1}_{\alpha}}}$) are both bounded operators on $L^{2}(\mathbb{T}^{2})$.
@@
Let $\phi$ be a rational inner function on $\mathbb{D}^{2}$ and let $\alpha\in\mathbb{T}$ be generic for $\phi$. Then $J_{\alpha}$ is surjective, $\vartheta^{1}_{\alpha}$ is bounded, and $U^{1}_{\alpha}=J_{\alpha}^{*}M_{z_{1}}J_{\alpha}$, where
@@
Since $\alpha$ is generic for $\phi$, Theorem 4.2 in [4] implies that $J_{\alpha}$ is unitary. Since $\phi$ is bounded on $\mathbb{D}^{2}$, so is $B_{1}\phi$. If the function
@@
were not bounded on $\mathbb{D}$, then there would exist a $\tau\in\mathbb{T}$ and a sequence $(z_{n})\subseteq\mathbb{D}$ converging to $\tau$ such that $\phi(0,z_{n})$ converged to $\alpha.$ Since $\phi$ does not have any singularities on $\mathbb{D}\times\mathbb{T}$, we can conclude that $\phi(0,\tau)=\alpha.$ Consider the finite Blaschke product $\phi_{\tau}(z_{1}):=\phi(z_{1},\tau)$. Since $\phi_{\tau}(0)=\alpha$, this means $\phi_{\tau}$ must be identically $\alpha$. Thus, the set $\mathbb{T}\times\{\tau\}$ is contained in the level set $\mathcal{C}_{\alpha}$. This implies that $\alpha$ is not generic, a contradiction. Thus, $\vartheta^{1}_{\alpha}$ must be bounded on $\mathbb{D}^{2}$ and the rest of the statement follows from Theorem 4.3. ∎
@@
By known properties of two-variable Schur functions (see for example [8]), there exist positive kernels $K_{1},K_{2}\colon\mathbb{D}^{2}\times\mathbb{D}^{2}\rightarrow\mathbb{C}$ such that for all $z,w\in\mathbb{D}^{2},$
$k^{\phi}_{w}(z)=\frac{1-\phi(z)\overline{\phi(w)}}{(1-z_{1}\bar{w}_{1})(1-z_{2}\bar{w}_{2})}=\frac{K_{1}(z,w)}{1-z_{1}\bar{w}_{1}}+\frac{K_{2}(z,w)}{1-z_{2}\bar{w}_{2}}.$ (4.5)
@@
$=\left\langle g_{w},\bar{\alpha}\frac{B_{1}\phi}{1-\bar{\alpha}\phi(0,\cdot)}k_{(0,\zeta)}\right\rangle_{L^{2}}$
@@
for some $c\neq 0.$ Since $k_{(0,w_{2})}\not\in K_{\phi},$ $P_{\phi}$ is required to ensure that the operator $U_{\alpha}^{1}$ maps into $K_{\phi}$. ∎
Remark.
It is worth noting that if $\phi$ is a rational inner function that both depends on $z_{1}$ and $z_{2}$ and does not have a factor of $z_{1}$, then we can apply Proposition 4.7 to $\phi$ for all of its generic $\alpha\in\mathbb{T}.$ In particular, as discussed in Corollary 4.6 and its proof, $J_{\alpha}$ is surjective and $\frac{1}{1-\bar{\alpha}\phi(0,z_{2})}$ is bounded. For a more specific (and simpler) example of a $U_{\alpha}^{1}$ requiring $P_{\phi}$, see Subsection 6.1.
@@ 5. RIF Level Sets and Taylor Joint Spectrum
We are interested in the joint spectrum of the shift operators $(M_{z_{1}},M_{z_{2}})$ on a given space of functions $L^{2}(\mu)$.
@@
Let $\mu$ be a finite positive Borel measure on $\mathbb{T}^{2}$ with no point-masses and let $M_{z_{1}}$, $M_{z_{2}}$ denote multiplication by $z_{1},z_{2}$ on $L^{2}(\mu)$. Assume that $\mu$ satisfies the following two conditions:
  • (a)
    For all $\lambda\in\mathbb{T},$ $\mu\left(\{\lambda\}\times\mathbb{T}\right)=0=\mu\left(\mathbb{T}\times\{\lambda\}\right),$
  • (b)
    There is a dense subset $\mathcal{S}$ of $\mathrm{supp}\,\mu$ with the following property: for each $(\lambda_{1},\lambda_{2})\in\mathcal{S}$, there is some $c>0$ such that for $\epsilon>0$ sufficiently small, $\mu(V_{\epsilon}(\lambda_{1},\lambda_{2}))\geq c\epsilon^{2}.$
Then $\boldsymbol{\sigma}(M_{z_{1}},M_{z_{2}})=\mathrm{supp}\,\mu.$
@@
It was pointed out to us by Michael Hartz that the same conclusion, namely that $\boldsymbol{\sigma}(M_{z_{1}},M_{z_{2}})=\mathrm{supp}\,\mu$, holds without the additional hypotheses (a) and (b), as can be seen by appealing for instance to a result of Vasilescu (see [38, Theorem 1.1]). We have opted to present a direct proof so as to keep our paper self-contained.
@@
For the first part of the proof, we will establish that $\boldsymbol{\sigma}(M_{z_{1}},M_{z_{2}})\subseteq\mathrm{supp}\,\mu$ by showing if $(\lambda_{1},\lambda_{2})\not\in\mathrm{supp}\,\mu$, then $(\lambda_{1},\lambda_{2})\not\in\boldsymbol{\sigma}(M_{z_{1}},M_{z_{2}})$. This means that we need to show that $\text{Ker}(\delta_{n})\subseteq\text{Ran}(\delta_{n-1})$ for $n=1,2,3$.
@@
$h(z_{1},z_{2})=\left\{\begin{array}[]{ll}\tfrac{f_{1}(z_{1},z_{2})}{z_{1}-\lambda_{1}}&\text{ for }(z_{1},z_{2})\in\mathbb{T}^{2}\setminus(\{\lambda_{1}\}\times\mathbb{T})\\ \tfrac{f_{2}(z_{1},z_{2})}{z_{2}-\lambda_{2}}&\text{ for }(z_{1},z_{2})\in\mathbb{T}^{2}\setminus(\mathbb{T}\times\{\lambda_{2}\})\end{array}\right$.
By assumption (a) and (5.1), $h$ is well-defined $\mu$ a.e. on $\mathbb{T}^{2}$. We just need to show that $h\in L^{2}(\mu)$. Then we will have
$\delta_{1}(h)=\left((M_{z_{1}}-\lambda_{1})h,(M_{z_{2}}-\lambda_{2})h\right)=\left(f_{1},f_{2}\right)$
and $(f_{1},f_{2})\in\text{Range}(\delta_{1})$, as needed. Because $(\lambda_{1},\lambda_{2})\not\in\mathrm{supp}\,\mu$, there is an $\epsilon>0$ so that $\mu(V_{\epsilon}(\lambda_{1},\lambda_{2}))=0$. To show $h\in L^{2}(\mu)$, define
@@
so $\mathbb{T}^{2}=S_{1}\cup S_{2}\cup V_{\epsilon}(\lambda_{1},\lambda_{2})$. Then we have
@@
$f_{2}=\frac{h}{z_{1}-\lambda_{1}}1_{S_{1}}\text{ and }f_{1}=\frac{h}{z_{2}-\lambda_{2}}1_{S_{2}},$
@@
for $\mu$ a.e. $(z_{1},z_{2})\in\mathbb{T}^{2}$. Then for $\epsilon>0$ sufficiently small,
@@
since the $\lim_{\epsilon\rightarrow 0}|h_{1}(z)1_{V_{\epsilon}(\lambda_{1},\lambda_{2})}(z)|^{2}=0=\lim_{\epsilon\rightarrow 0}|h_{2}(z)1_{V_{\epsilon}(\lambda_{1},\lambda_{2})}(z)|^{2}$ for $\mu$-a.e. $z\in\mathbb{T}^{2}$ (where we used the fact that $\mu$ has no point-masses).
@@
Let $\phi$ be a rational inner function on $\mathbb{D}^{2}$ and let $\alpha\in\mathbb{T}$ be generic for $\phi$. Fix $(\lambda_{1},\lambda_{2})$ in $\mathcal{C}_{\alpha}$ such that $(\lambda_{1},\lambda_{2})$ is not a singularity of $\phi$. Then, there is some $c>0$ such that for $\epsilon>0$ sufficiently small,
@@
We first need some preliminary facts. Note that since $\alpha$ is generic and $\phi(\lambda_{1},\lambda_{2})=\alpha$, then the one-variable functions $\phi(\cdot,\lambda_{2})$ and $\phi(\lambda_{1},\cdot)$ cannot be constant. Since non-constant finite Blaschke products have non-zero derivatives on $\mathbb{T}$, we have $\frac{\partial\phi}{\partial z_{j}}(\lambda_{1},\lambda_{2})\neq 0$ for $j=1,2$. By continuity, for $\epsilon>0$ sufficiently small, there is an $m>0$ such that $|\frac{\partial\phi}{\partial z_{1}}(z_{1},z_{2})|<m$ for all $(z_{1},z_{2})\in V_{\epsilon}(\lambda_{1},\lambda_{2})$.
Furthermore, the implicit function theorem implies that $\mathcal{C}_{\alpha}$ is parameterized by a single analytic function $z_{2}=g_{\alpha}(z_{1})$ in a neighborhood of $(\lambda_{1},\lambda_{2})$. It is also worth noting that $g_{\alpha}^{\prime}(\lambda_{1})\neq 0$. To see this, observe that
@@
$g_{\alpha}(z_{1})=\lambda_{2}+a(\lambda_{1}-z_{1})+O(|\lambda_{1}-z_{1}|^{2}).$ (5.3)
@@
$\geq\int_{\mathbb{T}}1_{V_{\frac{\epsilon}{2}}(\lambda_{1},\lambda_{2})}(\zeta,g_{\alpha}(\zeta))\frac{1}{|\frac{\partial\phi}{\partial z_{1}}(\zeta,g_{\alpha}(\zeta))|}dm(\zeta)$
@@
Setting $b=\min\{\frac{\epsilon}{4},b_{1}\}$, we can see that if $|\tau_{1}-\theta|<b\epsilon$, then
@@
Now we can compute the Taylor joint spectrum of the shifts $(M_{z_{1}},M_{z_{2}})$ on the Clark spaces associated to rational inner functions.
Corollary 5.4.
Let $\phi$ be a rational inner function on $\mathbb{D}^{2}$ and let $\alpha\in\mathbb{T}$ be generic for $\phi$. Then on $L^{2}(\sigma_{\alpha})$, we have $\boldsymbol{\sigma}(M_{z_{1}},M_{z_{2}})=\mathcal{C}_{\alpha}$.

### v2

@@ 3.4. Taylor Joint Spectrum
While there are several different notions of spectrum for tuples of commuting bounded operators, the Taylor joint spectrum has proven to be a particularly useful tool in the study of shift operators on Hilbert spaces of analytic functions; for instances of this, see [24, 29, 10] and their lists of references. Comprehensive treatments of the Taylor joint spectrum for tuples of operators can be found in [42, 25] or [36, Chapter IV]. These references also include an overview of other spectral notions (the Harte spectrum, etc.) and conditions for when they coincide. Since we deal exclusively with pairs of operators, we only provide the relevant definitions in that setting.
@@ 4. Clark Unitary Operators $(U_{\alpha}^{1},U_{\alpha}^{2})$
Let $\phi$ be a nonconstant inner function in $\mathbb{D}^{2}$. We begin by computing the adjoint of the associated Clark embedding $J_{\alpha}$ which was defined in Subsection 3.2. The following result is a direct analog of Clark’s formula for the one-variable adjoint, see e.g., [28, Theorem 11.6].
@@
(4.2) $\vartheta^{1}_{\alpha}(z):=\bar{\alpha}\frac{(B_{1}\phi)(z)}{1-\bar{\alpha}\phi(0,z_{2})}.$
@@
Now we will use Proposition 4.2 to connect $M_{z_{1}}$ on $L^{2}(\sigma_{\alpha})$ with the compressed shift $S_{\phi}^{1}$ on $K_{\phi}$ for certain $\phi$. To do that, recall that $H^{2}_{2}$ is the Hardy space on the disk with independent variable $z_{2}$. If $f\in H^{2}(\mathbb{D}^{2})$, then
@@
(4.3) $U^{1}_{\alpha}=S_{\phi}^{1}+P_{\phi}P_{H^{2}_{2}}M_{\overline{\vartheta^{1}_{\alpha}}}.$
@@
Let $U_{\alpha}^{1}=(V_{\alpha}^{1})^{*}$, where $V_{\alpha}^{1}$ was defined in Proposition 4.2. By Proposition 4.2 and the assumption that $J_{\alpha}$ is a surjective isometry (and hence unitary), we have $M_{z_{1}}=J_{\alpha}U^{1}_{\alpha}J^{*}_{\alpha}$ or equivalently, $U_{\alpha}^{1}=J_{\alpha}^{*}M_{z_{1}}J_{\alpha}$. Since $M_{z_{1}}$ is unitary on $L^{2}(\sigma_{\alpha})$, the operator $U_{\alpha}^{1}$ is unitary on $K_{\phi}$. Furthermore, since $\vartheta^{1}_{\alpha}$ is bounded, multiplication by $\vartheta^{1}_{\alpha}$ and by $\overline{\vartheta^{1}_{\alpha}}$ (denoted $M_{\vartheta^{1}_{\alpha}}$ and $M_{\overline{\vartheta^{1}_{\alpha}}}$) are both bounded operators on $L^{2}(\mathbb{T}^{2})$.
@@
Let $\phi$ be a rational inner function on $\mathbb{D}^{2}$ with $\deg_{j}\phi\geq 1$ for $j=1,2$ and let $\alpha\in\mathbb{T}$ be generic for $\phi$. Then $J_{\alpha}$ is surjective, $\vartheta^{1}_{\alpha}$ is bounded, and $U^{1}_{\alpha}=J_{\alpha}^{*}M_{z_{1}}J_{\alpha}$, where
@@
Since $\alpha$ is generic for $\phi$, Theorem 3.4 implies that $J_{\alpha}$ is unitary. Since $\phi$ is bounded on $\mathbb{D}^{2}$, so is $B_{1}\phi$. If the function
@@
were not bounded on $\mathbb{D}$, then there would exist a $\tau\in\mathbb{T}$ and a sequence $(z_{n})\subseteq\mathbb{D}$ converging to $\tau$ such that $\phi(0,z_{n})$ converged to $\alpha.$ Since $\phi$ does not have any singularities on $\mathbb{D}\times\mathbb{T}$, we can conclude that $\phi(0,\tau)=\alpha.$ Consider the finite Blaschke product $\phi_{\tau}(z_{1}):=\phi(z_{1},\tau)$. Since $\phi_{\tau}(0)=\alpha$, this means $\phi_{\tau}$ must be identically $\alpha$. Thus, the set $\mathbb{T}\times\{\tau\}$ is contained in the level set $\mathcal{C}_{\alpha}$. This implies that $\alpha$ is not generic for $\phi$, a contradiction. Thus, $\vartheta^{1}_{\alpha}$ must be bounded on $\mathbb{D}^{2}$ and the rest of the statement follows from Theorem 4.3. ∎
@@
As was explained in Subsection 3.3 (see also [13], for example), there exist positive kernels $K_{1},K_{2}\colon\mathbb{D}^{2}\times\mathbb{D}^{2}\rightarrow\mathbb{C}$ such that for all $z,w\in\mathbb{D}^{2},$
(4.5) $k^{\phi}_{w}(z)=\frac{1-\phi(z)\overline{\phi(w)}}{(1-z_{1}\bar{w}_{1})(1-z_{2}\bar{w}_{2})}=\frac{K_{1}(z,w)}{1-z_{1}\bar{w}_{1}}+\frac{K_{2}(z,w)}{1-z_{2}\bar{w}_{2}}.$
@@
$=\left\langle g_{w},\bar{\alpha}\frac{B_{1}\phi}{1-\bar{\alpha}\phi(0,\cdot)}k_{(0,\zeta)}\right\rangle_{H^{2}}$
$=\alpha\left\langle g_{w},T_{1}^{*}\left(\frac{\phi}{1-\bar{\alpha}\phi(0,\cdot)}k_{(0,\zeta)}\right)\right\rangle_{H^{2}}$
@@
for some $c\neq 0.$ Since $k_{(0,w_{2})}\not\in K_{\phi},$ the projection $P_{\phi}$ is required to ensure that the operator $U_{\alpha}^{1}$ maps into $K_{\phi}$. ∎
Remark 4.8.
It is worth noting that if $\phi$ is a rational inner function that both depends on $z_{1}$ and $z_{2}$ and does not have a factor of $z_{1}$, then we can apply Proposition 4.7 (and Theorem 4.9 below) to $\phi$ for all of its generic $\alpha\in\mathbb{T}.$ In particular, as discussed in Corollary 4.6 and its proof, $J_{\alpha}$ is surjective and $\frac{1}{1-\bar{\alpha}\phi(0,z_{2})}$ is bounded and hence in $H^{2}_{2}$. For more details about an example of a $U_{\alpha}^{1}$ requiring $P_{\phi}$, see Subsection 6.1.
While the perturbations needed to complete a compressed shift $S_{1}^{\phi}$ to a unitary are often infinite-rank operators, they are not too large in the following sense.
Theorem 4.9.
Let $J_{\alpha}:K_{\phi}\rightarrow L^{2}(\sigma_{\alpha})$ be surjective and let $\vartheta^{1}_{\alpha}$ be bounded on $\mathbb{D}^{2}$, with $\frac{1}{1-\bar{\alpha}\phi(0,z_{2})}\in H^{2}_{2}.$ As in (3.6), write $K_{\phi}=S_{1}^{\mathrm{max}}\oplus S^{\mathrm{min}}_{2}$. Then
$P_{\phi}P_{H^{2}_{2}}M_{\overline{\vartheta_{\alpha}}}\equiv 0\quad\textrm{on}\quad S_{1}^{\mathrm{max}}.$
Proof.
Fix $f\in S_{1}^{\mathrm{max}}$ and $w_{2}\in\mathbb{D}.$ By our assumptions
$g_{w_{2}}(z_{2}):=\frac{1}{1-\bar{\alpha}\phi(0,z_{2})}\frac{1}{1-\bar{w}_{2}z_{2}}\in H^{2}_{2}.$
Proposition 3.5 in [11], which says $(B_{1}\phi)H^{2}_{2}\subseteq S^{\mathrm{min}}_{2}$, implies that
$h_{w_{2}}(z):=\frac{\vartheta^{1}_{\alpha}(z)}{1-\bar{w}_{2}z_{2}}=\bar{\alpha}(B_{1}\phi)(z)g_{w_{2}}(z_{2})\in S^{\mathrm{min}}_{2}.$
From this, we can conclude that
$\left(P_{H^{2}_{2}}M_{\overline{\vartheta^{1}_{\alpha}}}f\right)(0,w_{2})$ $=\left\langle P_{H^{2}_{2}}M_{\overline{\vartheta^{1}_{\alpha}}}f,k_{(0,w_{2})}\right\rangle_{H^{2}_{2}}$
$=\left\langle M_{\overline{\vartheta^{1}_{\alpha}}}f,k_{(0,w_{2})}\right\rangle_{L^{2}}$
$=\left\langle f,h_{w_{2}}\right\rangle_{H^{2}}=0,$
since $S_{1}^{\mathrm{max}}\perp_{H^{2}}S^{\mathrm{min}}_{2}.$ This immediately implies that
$P_{\phi}P_{H^{2}_{2}}M_{\overline{\vartheta_{\alpha}}}f=P_{\phi}0=0,$
which completes the proof. ∎
@@ 5. RIF Level Sets and Taylor Joint Spectrum
We are interested in the Taylor joint spectrum of the shift operators $(M_{z_{1}},M_{z_{2}})$ on a given space of functions $L^{2}(\mu)$.
@@
Let $\mu$ be a finite positive Borel measure on $\mathbb{T}^{2}$ and assume that $\mu$ behaves like Lebesgue measure in the following sense: there is a dense subset $\mathcal{S}$ of $\mathrm{supp}\,\mu$ such that for each $(\lambda_{1},\lambda_{2})\in\mathcal{S}$, there is some $c>0$ such that for $\epsilon>0$ sufficiently small, $\mu(V_{\epsilon}(\lambda_{1},\lambda_{2}))\geq c\epsilon^{2}.$
Let $M_{z_{1}}$, $M_{z_{2}}$ denote multiplication by $z_{1},z_{2}$ on $L^{2}(\mu)$. Then $\boldsymbol{\sigma}(M_{z_{1}},M_{z_{2}})=\mathrm{supp}\,\mu.$
@@
It was pointed out to us by Michael Hartz that the same conclusion, namely that $\boldsymbol{\sigma}(M_{z_{1}},M_{z_{2}})=\mathrm{supp}\,\mu$, holds without the additional hypothesis on $\mu$, as can be seen by appealing for instance to a result of Vasilescu (see [44, Theorem 1.1]). We have opted to present a direct proof so as to keep our paper self-contained and highlight some specific properties of Clark measures associated to rational inner functions.
@@
For the first part of the proof, we will establish that $\boldsymbol{\sigma}(M_{z_{1}},M_{z_{2}})\subseteq\mathrm{supp}\,\mu$ by showing if $(\lambda_{1},\lambda_{2})\not\in\mathrm{supp}\,\mu$, then $(\lambda_{1},\lambda_{2})\not\in\boldsymbol{\sigma}(M_{z_{1}},M_{z_{2}})$. This means that we need to show that $\text{Ker}(\delta_{n})\subseteq\text{Ran}(\delta_{n-1})$ for $n=1,2,3$. (For the definitions of $\delta_{1},\delta_{2},\delta_{3},$ see Subsection 3.4.)
@@
$h(z_{1},z_{2})=\left\{\begin{array}[]{ll}\tfrac{f_{1}(z_{1},z_{2})}{z_{1}-\lambda_{1}}&\text{ for }(z_{1},z_{2})\in\mathbb{T}^{2}\setminus(\{\lambda_{1}\}\times\mathbb{T})\\ \tfrac{f_{2}(z_{1},z_{2})}{z_{2}-\lambda_{2}}&\text{ for }(z_{1},z_{2})\in\mathbb{T}^{2}\setminus(\mathbb{T}\times\{\lambda_{2}\}).\end{array}\right$.
By assumption (5.1) and since $(\lambda_{1},\lambda_{2})\not\in\mathrm{supp}\,\mu,$ $h$ is well-defined $\mu$-a.e. on $\mathbb{T}^{2}$. Furthermore, note that for $\mu$-a.e. $(z_{1},z_{2})\in\mathbb{T}^{2}\setminus(\{\lambda_{1}\}\times\mathbb{T})$, the definition of $h$ implies
$(M_{z_{1}}-\lambda_{1})h(z_{1},z_{2})=f_{1}(z_{1},z_{2}).$
Meanwhile, for $\mu$-a.e. $(z_{1},z_{2})\in\mathbb{T}^{2}\setminus(\mathbb{T}\times\{\lambda_{2}\})$, the definition of $h$ and (5.1) implies that
$(M_{z_{1}}-\lambda_{1})h(z_{1},z_{2})=\left(\frac{z_{1}-\lambda_{1}}{z_{2}-\lambda_{2}}\right)f_{2}(z_{1},z_{2})=f_{1}(z_{1},z_{2}).$
Similar arguments hold for $(M_{z_{2}}-\lambda_{2})h$ and so
(5.2) $\left((M_{z_{1}}-\lambda_{1})h,(M_{z_{2}}-\lambda_{2})h\right)=\left(f_{1},f_{2}\right)$
$\mu$-a.e. Now, we just need to show that $h\in L^{2}(\mu)$. Once that is established, (5.2) will give $\delta_{1}(h)=\left(f_{1},f_{2}\right)$, so $(f_{1},f_{2})\in\text{Range}(\delta_{1})$, as needed.
To show $h\in L^{2}(\mu)$, note that since $(\lambda_{1},\lambda_{2})\not\in\mathrm{supp}\,\mu$, there is an $\epsilon>0$ so that $\mu(V_{\epsilon}(\lambda_{1},\lambda_{2}))=0$. To show $h\in L^{2}(\mu)$, define
@@
so $\mathbb{T}^{2}=S_{1}\cup S_{2}\cup V_{\epsilon}(\lambda_{1},\lambda_{2})$. Then choosing $\epsilon$ sufficiently small gives
@@
$f_{2}=\frac{h}{z_{1}-\lambda_{1}}1_{S_{1}}\text{ and }f_{1}=\frac{-h}{z_{2}-\lambda_{2}}1_{S_{2}},$
@@
for $\mu$ a.e. $(z_{1},z_{2})\in\mathbb{T}^{2}$. First, if $\mu$ has a point-mass at $(\lambda_{1},\lambda_{2})$ with weight $c\neq 0$, then $h_{1},h_{2}$ would have to be well-defined at $(\lambda_{1},\lambda_{2})$ and so
$c=\mu(\{(\lambda_{1},\lambda_{2})\})=c(\lambda_{1}-\lambda_{1})h_{2}(\lambda_{1},\lambda_{2})+c(\lambda_{2}-\lambda_{2})h_{1}(\lambda_{1},\lambda_{2})=0,$
a contradiction.
Now assume that $\mu$ does not have a point-mass at $(\lambda_{1},\lambda_{2}).$ Then for $\epsilon>0$ sufficiently small,
@@
since the $\lim_{\epsilon\rightarrow 0}|h_{1}(z)1_{V_{\epsilon}(\lambda_{1},\lambda_{2})}(z)|^{2}=0=\lim_{\epsilon\rightarrow 0}|h_{2}(z)1_{V_{\epsilon}(\lambda_{1},\lambda_{2})}(z)|^{2}$ for $\mu$-a.e. $z\in\mathbb{T}^{2}$ (where we used the fact that $\mu$ does not have a point mass at $(\lambda_{1},\lambda_{2})$).
@@
Let $\phi=\frac{\tilde{p}}{p}$ be a bidegree $(n_{1},n_{2})$ RIF with $n_{1},n_{2}\geq 1$ and let $\alpha\in\mathbb{T}$ be generic for $\phi$. Fix $(\lambda_{1},\lambda_{2})$ in $\mathcal{C}_{\alpha}$ such that $(\lambda_{1},\lambda_{2})$ is not a singularity of $\phi$. Then, there is some $c>0$ such that for $\epsilon>0$ sufficiently small,
@@
We first need some preliminary facts. Note that since $\alpha$ is generic and $\phi(\lambda_{1},\lambda_{2})=\alpha$, then the one-variable functions $\phi(\cdot,\lambda_{2})$ and $\phi(\lambda_{1},\cdot)$ cannot be constant. Since non-constant finite Blaschke products have non-zero derivatives on $\mathbb{T}$, we have $\frac{\partial\phi}{\partial z_{j}}(\lambda_{1},\lambda_{2})\neq 0$ for $j=1,2$. By continuity, for $\epsilon>0$ sufficiently small, there is an $m>0$ such that $|\frac{\partial\phi}{\partial z_{2}}(z_{1},z_{2})|<m$ for all $(z_{1},z_{2})\in V_{\epsilon}(\lambda_{1},\lambda_{2})$.
As described in Subsection 3.2, $\mathcal{C}_{\alpha}$ is globally parameterized by a finite family of analytic functions on $\mathbb{T}^{2}$. However, the implicit function theorem implies that $\mathcal{C}_{\alpha}$ is actually parameterized by a single analytic function $z_{2}=g_{\alpha}(z_{1})$ in a neighborhood of $(\lambda_{1},\lambda_{2})$. It is also worth noting that $g_{\alpha}^{\prime}(\lambda_{1})\neq 0$. To see this, observe that
@@
(5.4) $g_{\alpha}(z_{1})=\lambda_{2}+a(\lambda_{1}-z_{1})+O(|\lambda_{1}-z_{1}|^{2}).$
@@
$\geq\int_{\mathbb{T}}1_{V_{\frac{\epsilon}{2}}(\lambda_{1},\lambda_{2})}(\zeta,g_{\alpha}(\zeta))\frac{1}{|\frac{\partial\phi}{\partial z_{2}}(\zeta,g_{\alpha}(\zeta))|}dm(\zeta)$
@@
Setting $b=\min\{\frac{1}{4},b_{1}\}$, we can see that if $|\tau_{1}-\theta|<b\epsilon$, then
@@
Theorem 5.4.
Let $\phi=\frac{\tilde{p}}{p}$ be a bidegree $(n_{1},n_{2})$ RIF with $n_{1},n_{2}\geq 1$ and let $\alpha\in\mathbb{T}$ be generic for $\phi$. Then there is a commuting pair of unitary perturbations $(U_{\alpha}^{1},U_{\alpha}^{2})$ of the compressed shifts $(S_{\phi}^{1},S_{\phi}^{2})$ on $K_{\phi}$ such that their Taylor joint spectrum $\boldsymbol{\sigma}(U_{\alpha}^{1},U_{\alpha}^{2})=\mathcal{C}_{\alpha}$.